OpenStudy (anonymous):

Which shows the values in increasing order? 1 over 5, 3 for every 8, 45%, 0.5 0.5, 1 over 5, 45%, 3 for every 8 1 over 5, 3 for every 8, 0.5, 45% 0.5, 1 over 5, 3 for every 8, 45% (This is answer is repeated in the answers section on the actual page.)

OpenStudy (asnaseer):

convert them all to decimal and then order them

OpenStudy (anonymous):

Well, the thing is, the lesson wasn't really converting so easily, so I don't get the conversion.

OpenStudy (asnaseer):

do you know what 1/5 is in decimal notation?

OpenStudy (anonymous):

0.1?

OpenStudy (asnaseer):

\[\frac{1}{5}=\frac{2}{10}=0.2\]

OpenStudy (anonymous):

._. oh....

OpenStudy (asnaseer):

what about percentages - do you know what 45% is in decimal?

OpenStudy (anonymous):

0.5 > 0.45 > 3 for every 8 (3/8=0,375) > 1 over 5 (0.2)

OpenStudy (anonymous):

0.45? or 0.045?

OpenStudy (anonymous):

%45 = 0,45 3 of every 8 = 3/8 = 0,375 1 over 5 = 1/5 = 0,2

OpenStudy (asnaseer):

\[45\text{%}=\frac{45}{100}=0.45\]

OpenStudy (asnaseer):

remember % means "out of a 100"

OpenStudy (anonymous):

ok ._.

OpenStudy (asnaseer):

so that just leaves 3/8

OpenStudy (asnaseer):

are you allowed to use a calculator?

OpenStudy (anonymous):

yes._.

OpenStudy (asnaseer):

ok - so you should be able to work out 3/8 on a calculator - what do you get for that?

OpenStudy (anonymous):

Wouldn't that leave me with 10?

OpenStudy (asnaseer):

how? you need to divide 3 by 8 on a calculator

OpenStudy (anonymous):

oh... I was thinking like this 3 x 2 = 6 8 x 2 = 16 6/16 = 6 is 10 away from 16.....

OpenStudy (anonymous):

3 divided by 8 = .375

OpenStudy (asnaseer):

correct, so now you have all the values in decimal: 1 over 5 = 1/5 = 0.2 3 for every 8 = 3/8 = 0.375 45% = 45/100 = 0.45 0.5 so you just need to order the following numbers: 0.2, 0.375, 0.45, 0.5

OpenStudy (anonymous):

thank you but, do you mind taking a look at the answers in the question again? The last answer is a repeat of the second one....even o nthe test i'm taking....

OpenStudy (anonymous):

wait, wait. no, no it is not, my bad.

OpenStudy (asnaseer):

:)

OpenStudy (anonymous):

just looked at the real test

OpenStudy (asnaseer):

do you understand how to order the resulting decimal numbers now?

OpenStudy (anonymous):

yes, thank you :)

OpenStudy (asnaseer):

yw - I'm glad I could help.